OpenStudy (ltrout):

[Fan + Medal] Find the real or imaginary solution of the equation below by factoring. 4x^2 = -4x -1.

OpenStudy (caozeyuan):

\[4x ^{2}+4x=-1\]\[x(x+1)=-\frac{ 1 }{ 4 }\]

OpenStudy (ltrout):

@caozeyuan Wouldn't that only be x^2 + x = -1/4?

OpenStudy (caozeyuan):

I skiped a step, if you divide LHS and RHS of top eqn by 4, you get the eqn you just wrote

OpenStudy (ltrout):

Ah. Makes sense now. Could you help on one more? I'm not the best at factoring ¯\_(ツ)_/¯

OpenStudy (caozeyuan):

sure

OpenStudy (ltrout):

Okay, it's: 2x^2 + 3 = 4x

OpenStudy (ltrout):

Would I have to get the X's on one side and the 3 on the other side first?

OpenStudy (caozeyuan):

let me see, give me one sec

OpenStudy (caozeyuan):

divide by 2 to make the coeff of x^2 1 and then move all x's to LHS, and constants to RHS

OpenStudy (caozeyuan):

since b^2-4ac is less than 0 we know it has complex roots so we can assume x = a+bi, pultiply through and let imarginary part be 0 and real part be your constant

OpenStudy (ltrout):

So what would the final answer come out to be then? This is confusing.

OpenStudy (caozeyuan):

did you get the eqantion to be x(x-2)=-3/2

OpenStudy (ltrout):

Okay so it would be 2x^2 + 3 = 4x 2x^2 - 4x = -3. /2 x^2 - 2x = -3/2 x(x -2) = -3/2?

OpenStudy (caozeyuan):

yes, great

OpenStudy (caozeyuan):

do you know why it has no real roots?

OpenStudy (ltrout):

No, not really?

OpenStudy (caozeyuan):

a quadratic equation takes the form ax^2+bx+c=0 where a is not zero, correct?

OpenStudy (caozeyuan):

the determinant of this equation is b^2-4ac, and if it is negative we have complex roots only, and the roots are conjugate pair i.e. one is a+bi and the other is a-bi where i^2=-1

OpenStudy (ltrout):

Okay, that makes a lot of sense now. Thank you :)

OpenStudy (caozeyuan):

now, for this equation the determinant is negative so we dont have real roots

OpenStudy (caozeyuan):

therefore we can assume x=a+bi

OpenStudy (ltrout):

Are you good with polynomial long division? I need to check this answer I got.

OpenStudy (caozeyuan):

Not really, but I can try, do you still want me to solve the equation?

OpenStudy (ltrout):

No, that's fine. I should be able to get it from there. And I'm supposed to divide (x^3 + 7x^2 + 15x + 9) by (x+1) and the answer I got is x^2+6x+9.

OpenStudy (caozeyuan):

let me try

OpenStudy (caozeyuan):

yep, thats correct

OpenStudy (ltrout):

Alright, great! Thank you for helping me :)